OpenStudy (anonymous):

I'm trying to solve 2x^2+11x-21 by quadratic formula. Factoring gives 3/2, -7, but by the formula (which should always work), I get: (-11/4)+/-((22/4)(42)^(1/2)). Can anyone help me spot my stupid error?

OpenStudy (lalaly):

are u still there?

OpenStudy (anonymous):

factoring gives 3/2, -7, the quadratic formula is x=(-b±sqrt(b^2-4ac))/2a from your equation a=2,b=11,c=-21 put values in formula x=(-11±sqrt(121-(168)))/2x2 x=(-11±sqrt(289))/4 x=(-11±17)/4. x=3/2, -7

jhonyy9 (jhonyy9):

so if factoring this is right (2x-3)(x+7)=2x2+11x-21 so than your got solutions x=3/2 and x=-7 is right,correct sure always

jhonyy9 (jhonyy9):

mominanisar ATTENTION in this calcules are some errore because inside squarroot there not is minus 168 correct,right is plus 168 and in thgis way will get these correct factorized resultes

jhonyy9 (jhonyy9):

check it newly please

OpenStudy (anonymous):

i am seriously sorry actually this step is likex=(-11±sqrt(121-(168)))/2x2 new onex=(-11±sqrt(121-(-168)))/2x2 because 'c' is minus

OpenStudy (anonymous):

it is a typing mistake by me

OpenStudy (anonymous):

Thanks!